a man has to go 50m due north,40m due east and 20m due south to reach a field .The distance he has to walk to reach the field and the magnitude of displacement from his house to the field respectively are

Grade:9

2 Answers

Hasan Naqvi
8 years ago
Distance = 50 + 40 + 20 = 110m
Displacement= (302 + 402)1/2m  = 25001/2 m =50m
 
Jaswanth Sai Kandukuri
8 years ago
Magnitude of the displacement = (302+402)1/2 = (900+1600)1/2 ​= (2500)1/2 = 50